# VersaTune: Harnessing Vertical Domain Insights for Multi-Ability LLM Supervised Fine-Tuning

### v1

@@ Appendix C Evaluation Details
Appendix D More Experiment Results
D.1 Performance on Domain-Specific Tasks in Multi-Ability Fostering Scenario
We present herein the outcomes of the ablation study along with the raw scores from domain-specific benchmarks.
| Model | Method | Law | Medical | Finance | Science | Code | General | Avg. |
| LLaMA-2-7B | VersaTune Constant | 28.51↓ | 33.45↓ | 21.09↓ | 47.89↓ | 13.47↓ | 47.10↓ | 31.92↓ |
| VersaTune | 29.75 | 37.90 | 32.66 | 53.41 | 15.65 | 48.22 | 36.27 | |
| LLaMA-3-8B | VersaTune Constant | 42.80↓ | 33.74↓ | 31.60↓ | 55.41↓ | 31.55↓ | 49.48↓ | 40.76↓ |
| VersaTune | 43.77 | 43.97 | 41.19 | 58.64 | 32.28 | 51.20 | 45.18 | |
| Qwen-1.5-7B | VersaTune Constant | 36.75↓ | 27.94↓ | 46.29↓ | 47.86↓ | 34.04↓ | 56.59↓ | 41.58↓ |
| VersaTune | 38.00 | 31.36 | 56.08 | 54.44 | 36.15 | 57.24 | 45.55 | |
| Qwen-2-7B | VersaTune Constant | 39.29↓ | 40.01↓ | 50.73↓ | 56.61↓ | 51.29↓ | 60.57↓ | 49.75↓ |
| VersaTune | 43.05 | 43.36 | 64.67 | 61.00 | 52.90 | 60.85 | 54.31 | |
| Qwen-2.5-7B | VersaTune Constant | 41.99↓ | 39.96↓ | 55.71↓ | 64.00↓ | 55.63↓ | 60.82↓ | 53.02↓ |
| VersaTune | 44.20 | 43.19 | 65.79 | 71.15 | 60.97 | 61.82 | 57.85 | |
| LLaMA-2-13B | VersaTune Constant | 46.26↓ | 41.72↓ | 38.90↓ | 56.12↓ | 42.75↓ | 56.09↓ | 46.97↓ |
| VersaTune | 48.01 | 46.73 | 47.30 | 63.36 | 46.15 | 57.14 | 51.45 | |
| Qwen-2.5-14B | VersaTune Constant | 51.29↓ | 44.88↓ | 34.14↓ | 63.39↓ | 78.53↓ | 65.75↓ | 56.33↓ |
| VersaTune | 53.59 | 47.70 | 72.26 | 72.88 | 82.06 | 65.75 | 65.71 |
| Model | Method | LegalBench | LawBench |
| LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 |
| Inverse Distribution | 13.23↓ | 26.94↓ | |
| VersaTune Constant | 21.47↑ | 35.55↑ | |
| VersaTune | 23.18↑ | 36.31↑ | |
| LLaMA-3-8B | Uniform Distribution | 33.52 | 31.16 |
| Inverse Distribution | 27.83↓ | 27.48↓ | |
| VersaTune Constant | 47.85↑ | 37.75↑ | |
| VersaTune | 49.67↑ | 37.87↑ | |
| Qwen-1.5-7B | Uniform Distribution | 30.33 | 33.49 |
| Inverse Distribution | 29.80↓ | 30.13↓ | |
| VersaTune Constant | 37.97↑ | 35.52↑ | |
| VersaTune | 40.04↑ | 35.95↑ | |
| Qwen-2-7B | Uniform Distribution | 39.05 | 31.99 |
| Inverse Distribution | 34.01↓ | 27.81↓ | |
| VersaTune Constant | 45.86↑ | 32.72↑ | |
| VersaTune | 50.56↑ | 35.54↑ | |
| Qwen-2.5-7B | Uniform Distribution | 40.11 | 31.48 |
| Inverse Distribution | 36.36↓ | 26.98↓ | |
| VersaTune Constant | 48.78↑ | 35.20↑ | |
| VersaTune | 51.65↑ | 36.75↑ | |
| LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 |
| Inverse Distribution | 40.12↓ | 30.67↓ | |
| VersaTune Constant | 53.79↑ | 38.73↑ | |
| VersaTune | 55.87↑ | 40.14↑ | |
| Qwen-2.5-14B | Uniform Distribution | 50.73 | 39.49 |
| Inverse Distribution | 46.08↓ | 35.36↓ | |
| VersaTune Constant | 56.94↑ | 45.64↑ | |
| VersaTune | 60.59↑ | 46.58↑ |
| Model | Method | MedQA | MedMCQA |
| LLaMA-2-7B | Uniform Distribution | 23.45 | 27.57 |
| Inverse Distribution | 21.38↓ | 26.52↓ | |
| VersaTune Constant | 30.17↑ | 36.72↑ | |
| VersaTune | 35.04↑ | 40.75↑ | |
| LLaMA-3-8B | Uniform Distribution | 31.03 | 10.26 |
| Inverse Distribution | 25.51↓ | 8.77↓ | |
| VersaTune Constant | 37.33↑ | 30.15↑ | |
| VersaTune | 42.21↑ | 45.72↑ | |
| Qwen-1.5-7B | Uniform Distribution | 21.27 | 12.14 |
| Inverse Distribution | 16.15↓ | 9.91↓ | |
| VersaTune Constant | 25.60↑ | 30.28↑ | |
| VersaTune | 30.47↑ | 32.24↑ | |
| Qwen-2-7B | Uniform Distribution | 35.07 | 17.73 |
| Inverse Distribution | 23.90↓ | 16.31↓ | |
| VersaTune Constant | 40.89↑ | 39.13↑ | |
| VersaTune | 45.48↑ | 41.24↑ | |
| Qwen-2.5-7B | Uniform Distribution | 25.17 | 25.84 |
| Inverse Distribution | 24.16↓ | 19.35↓ | |
| VersaTune Constant | 30.20↑ | 49.71↑ | |
| VersaTune | 34.28↑ | 52.09↑ | |
| LLaMA-2-13B | Uniform Distribution | 32.98 | 36.54 |
| Inverse Distribution | 26.27↓ | 28.78↓ | |
| VersaTune Constant | 40.69↑ | 42.74↑ | |
| VersaTune | 45.78↑ | 47.67↑ | |
| Qwen-2.5-14B | Uniform Distribution | 47.85 | 38.71 |
| Inverse Distribution | 45.75↓ | 32.56↓ | |
| VersaTune Constant | 48.11↑ | 41.64↑ | |
| VersaTune | 50.24↑ | 45.15↑ |
| Model | Method | FinEval | FinanceIQ |
| LLaMA-2-7B | Uniform Distribution | 33.50 | 2.71 |
| Inverse Distribution | 32.96↓ | 2.53↓ | |
| VersaTune Constant | 35.89↑ | 6.28↑ | |
| VersaTune | 36.27↑ | 29.04↑ | |
| LLaMA-3-8B | Uniform Distribution | 34.83 | 4.97 |
| Inverse Distribution | 33.71↓ | 3.31↓ | |
| VersaTune Constant | 37.93↑ | 25.27↑ | |
| VersaTune | 38.80↑ | 43.58↑ | |
| Qwen-1.5-7B | Uniform Distribution | 53.87 | 14.17 |
| Inverse Distribution | 53.80↓ | 12.44↓ | |
| VersaTune Constant | 55.06↑ | 37.52↑ | |
| VersaTune | 55.52↑ | 56.63↑ | |
| Qwen-2-7B | Uniform Distribution | 59.49 | 14.62 |
| Inverse Distribution | 56.53↓ | 11.30↓ | |
| VersaTune Constant | 60.63↑ | 40.82↑ | |
| VersaTune | 60.95↑ | 68.39↑ | |
| Qwen-2.5-7B | Uniform Distribution | 59.58 | 31.66 |
| Inverse Distribution | 57.07↓ | 29.25↓ | |
| VersaTune Constant | 62.94↑ | 48.47↑ | |
| VersaTune | 62.48↑ | 69.09↑ | |
| LLaMA-2-13B | Uniform Distribution | 37.54 | 32.85 |
| Inverse Distribution | 36.67↓ | 26.76↓ | |
| VersaTune Constant | 39.33↑ | 38.47↑ | |
| VersaTune | 39.48↑ | 55.12↑ | |
| Qwen-2.5-14B | Uniform Distribution | 64.72 | 64.39 |
| Inverse Distribution | 64.88↑ | 60.53↓ | |
| VersaTune Constant | 65.03↑ | 73.24↑ | |
| VersaTune | 65.84↑ | 78.68↑ |
| Model | Method | SciEval | MMLU-Sci |
| LLaMA-2-7B | Uniform Distribution | 9.30 | 42.89 |
| Inverse Distribution | 8.98↓ | 39.67↓ | |
| VersaTune Constant | 49.91↑ | 45.87↑ | |
| VersaTune | 56.75↑ | 50.06↑ | |
| LLaMA-3-8B | Uniform Distribution | 6.51 | 50.17 |
| Inverse Distribution | 6.09↓ | 46.62↓ | |
| VersaTune Constant | 54.77↑ | 56.04↑ | |
| VersaTune | 56.67↑ | 60.61↑ | |
| Qwen-1.5-7B | Uniform Distribution | 7.63 | 50.67 |
| Inverse Distribution | 6.54↓ | 48.52↓ | |
| VersaTune Constant | 38.39↑ | 57.33↑ | |
| VersaTune | 47.76↑ | 61.11↑ | |
| Qwen-2-7B | Uniform Distribution | 25.30 | 62.73 |
| Inverse Distribution | 18.57↓ | 58.25↓ | |
| VersaTune Constant | 45.93↑ | 67.29↑ | |
| VersaTune | 51.58↑ | 70.42↑ | |
| Qwen-2.5-7B | Uniform Distribution | 19.88 | 65.84 |
| Inverse Distribution | 16.68↓ | 62.78↓ | |
| VersaTune Constant | 56.04↑ | 71.96↑ | |
| VersaTune | 68.14↑ | 74.16↑ | |
| LLaMA-2-13B | Uniform Distribution | 45.72 | 50.77 |
| Inverse Distribution | 38.96↓ | 48.68↓ | |
| VersaTune Constant | 57.13↑ | 55.10↑ | |
| VersaTune | 63.87↑ | 62.84↑ | |
| Qwen-2.5-14B | Uniform Distribution | 39.74 | 73.45 |
| Inverse Distribution | 27.68↓ | 68.22↓ | |
| VersaTune Constant | 48.31↑ | 78.46↑ | |
| VersaTune | 62.89↑ | 82.86↑ |
| Model | Method | HumanEval | MBPP |
| LLaMA-2-7B | Uniform Distribution | 5.67 | 3.44 |
| Inverse Distribution | 3.47↓ | 2.42↓ | |
| VersaTune Constant | 12.47↑ | 14.47↑ | |
| VersaTune | 15.62↑ | 15.68↑ | |
| LLaMA-3-8B | Uniform Distribution | 22.94 | 28.85 |
| Inverse Distribution | 19.67↓ | 24.34↓ | |
| VersaTune Constant | 29.88↑ | 33.22↑ | |
| VersaTune | 28.91↑ | 35.65↑ | |
| Qwen-1.5-7B | Uniform Distribution | 32.86 | 25.86 |
| Inverse Distribution | 28.55↓ | 23.94↓ | |
| VersaTune Constant | 38.41↑ | 29.67↑ | |
| VersaTune | 40.49↑ | 31.80↑ | |
| Qwen-2-7B | Uniform Distribution | 53.26 | 37.82 |
| Inverse Distribution | 50.65↓ | 33.63↓ | |
| VersaTune Constant | 56.71↑ | 45.87↑ | |
| VersaTune | 58.15↑ | 47.64↑ | |
| Qwen-2.5-7B | Uniform Distribution | 55.64 | 46.86 |
| Inverse Distribution | 52.97↓ | 44.63↓ | |
| VersaTune Constant | 59.15↑ | 52.10↑ | |
| VersaTune | 60.68↑ | 61.25↑ | |
| LLaMA-2-13B | Uniform Distribution | 36.54 | 38.55 |
| Inverse Distribution | 28.78↓ | 35.83↓ | |
| VersaTune Constant | 42.74↑ | 42.76↑ | |
| VersaTune | 47.67↑ | 44.62↑ | |
| Qwen-2.5-14B | Uniform Distribution | 68.75 | 72.14 |
| Inverse Distribution | 63.36↓ | 68.49↓ | |
| VersaTune Constant | 78.72↑ | 78.33↑ | |
| VersaTune | 82.64↑ | 81.48↑ |
| Model | Method | AGIEval | HellaSwag |
| LLaMA-2-7B | Uniform Distribution | 20.16 | 71.40 |
| Inverse Distribution | 18.83↓ | 71.33↓ | |
| VersaTune Constant | 22.31↑ | 71.89↑ | |
| VersaTune | 24.67↑ | 71.76↑ | |
| LLaMA-3-8B | Uniform Distribution | 23.87 | 73.26 |
| Inverse Distribution | 23.45↓ | 72.40↓ | |
| VersaTune Constant | 25.62↑ | 73.33↑ | |
[…]

### v3

@@ 5. Experiments
In this section, we first describe the details of our experimental setup (Section 5.1), including training datasets, models used for experimentation and benchmarks. We then outline the baseline methods we use for comparison (Section 5.2) and our experimental results to prove the effectiveness of VersaTune (Section 5.3).
| Model | Method | Law | Medical | Finance | Science | Code | General | ||||||
| LegalBench | LawBench | MedQA | MedMCQA | FinEval | FinanceIQ | SciEval | MMLU-Sci | HumanEval | MBPP | AGIEval | HellaSwag | ||
| LLaMA-2-7B | Uniform Distribution | 15.71 | 30.72 | 23.45 | 27.57 | 33.50 | 2.71 | 9.30 | 42.89 | 5.67 | 3.44 | 20.16 | 71.40 |
| Inverse Distribution | 13.23↓ | 26.94↓ | 21.38↓ | 26.52↓ | 32.96↓ | 2.53↓ | 8.98↓ | 39.67↓ | 3.47↓ | 2.42↓ | 18.83↓ | 71.33↓ | |
| VersaTune Constant | 21.47↑ | 35.55↑ | 30.17↑ | 36.72↑ | 35.89↑ | 6.28↑ | 49.91↑ | 45.87↑ | 12.47↑ | 14.47↑ | 22.31↑ | 71.89↑ | |
| VersaTune | 23.18↑ | 36.31↑ | 35.04↑ | 40.75↑ | 36.27↑ | 29.04↑ | 56.75↑ | 50.06↑ | 15.62↑ | 15.68↑ | 24.67↑ | 71.76↑ | |
| LLaMA-3-8B | Uniform Distribution | 33.52 | 31.16 | 31.03 | 10.26 | 34.83 | 4.97 | 6.51 | 50.17 | 22.94 | 28.85 | 23.87 | 73.26 |
| Inverse Distribution | 27.83↓ | 27.48↓ | 25.51↓ | 8.77↓ | 33.71↓ | 3.31↓ | 6.09↓ | 46.62↓ | 19.67↓ | 24.34↓ | 23.45↓ | 72.40↓ | |
| VersaTune Constant | 47.85↑ | 37.75↑ | 37.33↑ | 30.15↑ | 37.93↑ | 25.27↑ | 54.77↑ | 56.04↑ | 29.88↑ | 33.22↑ | 25.62↑ | 73.33↑ | |
| VersaTune | 49.67↑ | 37.87↑ | 42.21↑ | 45.72↑ | 38.80↑ | 43.58↑ | 56.67↑ | 60.61↑ | 28.91↑ | 35.65↑ | 28.78↑ | 73.62↑ | |
| Qwen-2-7B | Uniform Distribution | 39.05 | 31.99 | 35.07 | 17.73 | 59.49 | 14.62 | 25.30 | 62.73 | 53.26 | 37.82 | 47.31 | 73.60 |
| Inverse Distribution | 34.01↓ | 27.81↓ | 23.90↓ | 16.31↓ | 56.53↓ | 11.30↓ | 18.57↓ | 58.25↓ | 50.65↓ | 33.63↓ | 45.74↓ | 73.52↓ | |
| VersaTune Constant | 45.86↑ | 32.72↑ | 40.89↑ | 39.13↑ | 60.63↑ | 40.82↑ | 45.93↑ | 67.29↑ | 56.71↑ | 45.87↑ | 48.16↑ | 72.98↓ | |
| VersaTune | 50.56↑ | 35.54↑ | 45.48↑ | 41.24↑ | 60.95↑ | 68.39↑ | 51.58↑ | 70.42↑ | 58.15↑ | 47.64↑ | 48.02↑ | 73.67↑ | |
| Qwen-2.5-7B | Uniform Distribution | 40.11 | 31.48 | 25.17 | 25.84 | 59.58 | 31.66 | 19.88 | 65.84 | 55.64 | 46.86 | 45.42 | 73.69 |
| Inverse Distribution | 36.36↓ | 26.98↓ | 24.16↓ | 19.35↓ | 57.07↓ | 29.25↓ | 16.68↓ | 62.78↓ | 52.97↓ | 44.63↓ | 45.67↑ | 72.92↓ | |
| VersaTune Constant | 48.78↑ | 35.20↑ | 30.20↑ | 49.71↑ | 62.94↑ | 48.47↑ | 56.04↑ | 71.96↑ | 59.15↑ | 52.10↑ | 47.75↑ | 73.88↑ | |
| VersaTune | 51.65↑ | 36.75↑ | 34.28↑ | 52.09↑ | 62.48↑ | 69.09↑ | 68.14↑ | 74.16↑ | 60.68↑ | 61.25↑ | 49.73↑ | 73.90↑ | |
| LLaMA-2-13B | Uniform Distribution | 47.66 | 34.85 | 32.98 | 36.54 | 37.54 | 32.85 | 45.72 | 50.77 | 36.54 | 38.55 | 36.89 | 73.50 |
| Inverse Distribution | 40.12↓ | 30.67↓ | 26.27↓ | 28.78↓ | 36.67↓ | 26.76↓ | 38.96↓ | 48.68↓ | 28.78↓ | 35.83↓ | 36.67↓ | 73.11↓ | |
| VersaTune Constant | 53.79↑ | 38.73↑ | 40.69↑ | 42.74↑ | 39.33↑ | 38.47↑ | 57.13↑ | 55.10↑ | 42.74↑ | 42.76↑ | 37.91↑ | 74.27↑ | |
| VersaTune | 55.87↑ | 40.14↑ | 45.78↑ | 47.67↑ | 39.48↑ | 55.12↑ | 63.87↑ | 62.84↑ | 47.67↑ | 44.62↑ | 39.64↑ | 74.63↑ | |
| Qwen-2.5-14B | Uniform Distribution | 50.73 | 39.49 | 47.85 | 38.71 | 64.72 | 64.39 | 39.74 | 73.45 | 68.75 | 72.14 | 54.92 | 75.88 |
| Inverse Distribution | 46.08↓ | 35.36↓ | 45.75↓ | 32.56↓ | 64.88↑ | 60.53↓ | 27.68↓ | 68.22↓ | 63.36↓ | 68.49↓ | 54.87↓ | 75.42↓ | |
| VersaTune Constant | 56.94↑ | 45.64↑ | 48.11↑ | 41.64↑ | 65.03↑ | 73.24↑ | 48.31↑ | 78.46↑ | 78.72↑ | 78.33↑ | 55.04↑ | 76.45↑ | |
| VersaTune | 60.59↑ | 46.58↑ | 50.24↑ | 45.15↑ | 65.84↑ | 78.68↑ | 62.89↑ | 82.86↑ | 82.64↑ | 81.48↑ | 55.52↑ | 75.98↑ | |
5.1. Experimental Setup
Training Datasets. We have collected datasets spanning 6 domains for supervised fine-tuning (SFT), including Lawyer-Instruct333https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA (Jin et al., 2020), Finance Alpaca444https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations555https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca666https://github.com/sahil280114/codealpaca and Alpaca (Taori et al., 2023), denoted as , to represent SFT datasets with respect to law, medicine, finance, science, code as well as general capabilities. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other specific five domains, keeping only the general domain instances unrelated to them. Details of the training datasets can be found in Table 2.
Model Details. We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series as our pretrained language models . For scenarios aimed at fostering multi-ability, we trained and assessed models including LLaMA-2-7B, LLaMA-3-8B, LLaMA-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. In the context of domain expansion, the training and evaluations were performed using the LLaMA-3-8B and LLaMA-2-13B models. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for the Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Training Details. All experiments were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch, with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800.
Benchmarks. We evaluate the performance of the models on downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 3. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset (Zhang et al., 2024) for testing, aiming to ensure a robust and thorough evaluation.
5.2. Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. (3) VersaTune constant is implemented with fixed domain weights derived from the knowledge distribution obtained from probing the target model prior to fine-tuning, where we ablate the components of dynamic adaptation in Algorithm 2 for an in-depth analysis. Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
5.3. Results
We conduct evaluations to validate the efficiency of VersaTune across different open-source models in scenarios that encompass both effective multi-ability fostering and flexible domain expansion. We summarize the observations below.
| Target Domain | Training Step | Method | Variations in Comprehensive Domains () | Sum. () | ||||||
| (Epoch) | Law | Medicine | Finance | Science | Code | Other | Target | Non-Target | ||
| Law | 100% Specific Domain | 5.89 | 18.82 | 14.71 | 11.76 | 11.18 | 5.00 | 5.89 | 32.05 | |
| 1 | Uniform Distribution of Non-Target Domains | 5.38 | 7.35 | 6.47 | 6.58 | 0.05 | 13.99 | 5.38 | 6.48 | |
| VersaTune | 8.25 | 6.18 | 12.06 | 7.65 | 6.22 | 17.59 | 8.25 | 21.96 | ||
| 100% Specific Domain | 35.51 | 12.65 | 30.59 | 4.41 | 5.29 | 11.76 | 35.51 | 3.52 | ||
| 2 | Uniform Distribution of Non-Target Domains | 33.89 | 7.94 | 26.47 | 12.65 | 7.65 | 5.00 | 33.89 | 3.23 | |
| VersaTune | 35.14 | 3.53 | 18.82 | 6.89 | 17.06 | 13.24 | 35.14 | 45.76 | ||
| 100% Specific Domain | 55.84 | 17.94 | 12.35 | 8.82 | 23.53 | 5.00 | 55.84 | 42.94 | ||
| 3 | Uniform Distribution of Non-Target Domains | 52.89 | 8.24 | 15.00 | 12.06 | 10.17 | 5.12 | 52.89 | 20.59 | |
| VersaTune | 51.71 | 4.41 | 24.71 | 5.29 | 4.87 | 8.42 | 51.71 | 29.30 | ||
| 100% Specific Domain | 62.76 | 5.00 | 2.06 | 31.18 | 21.76 | 12.65 | 62.76 | 72.65 | ||
| 4 | Uniform Distribution of Non-Target Domains | 58.12 | 9.71 | 5.59 | 9.41 | 10.14 | 12.05 | 58.12 | 35.72 | |
| VersaTune | 59.08 | 5.59 | 13.82 | 8.82 | 5.61 | 6.17 | 59.08 | 12.37 | ||
| Medicine | 100% Specific Domain | 10.29 | 5.87 | 3.82 | 24.12 | 19.41 | 7.35 | 5.87 | 16.75 | |
| 1 | Uniform Distribution of Non-Target Domains | 3.82 | 5.36 | 5.29 | 5.59 | 2.65 | 6.53 | 5.36 | 0.36 | |
| VersaTune | 3.53 | 8.17 | 7.65 | 8.82 | 16.47 | 6.17 | 8.17 | 15.00 | ||
| 100% Specific Domain | 18.82 | 40.44 | 6.47 | 40.00 | 7.94 | 17.65 | 40.44 | 10.88 | ||
| 2 | Uniform Distribution of Non-Target Domains | 7.94 | 33.68 | 9.12 | 20.59 | 4.98 | 5.64 | 33.68 | 2.87 | |
| VersaTune | 12.35 | 35.21 | 7.65 | 8.84 | 16.52 | 12.65 | 35.21 | 58.01 | ||
| 100% Specific Domain | 22.35 | 58.78 | 8.82 | 7.94 | 19.12 | 10.00 | 58.78 | 52.35 | ||
| 3 | Uniform Distribution of Non-Target Domains | 10.46 | 55.69 | 14.98 | 12.35 | 4.41 | 14.18 | 55.69 | 31.68 | |
| VersaTune | 3.53 | 53.85 | 4.52 | 17.06 | 7.64 | 6.03 | 53.85 | 22.68 | ||
| 100% Specific Domain | 27.94 | 64.61 | 11.76 | 2.35 | 21.76 | 12.65 | 64.61 | 76.46 | ||
| 4 | Uniform Distribution of Non-Target Domains | 13.82 | 58.07 | 11.18 | 2.35 | 6.47 | 18.53 | 58.07 | 47.65 | |
| VersaTune | 5.59 | 59.81 | 4.27 | 10.68 | 5.94 | 10.26 | 59.81 | 15.38 | ||
| Finance | 100% Specific Domain | 20.59 | 7.94 | 5.45 | 10.29 | 12.65 | 6.47 | 5.45 | 16.76 | |
| 1 | Uniform Distribution of Non-Target Domains | 12.05 | 7.36 | 8.21 | 6.74 | 19.41 | 8.53 | 8.21 | 8.83 | |
| VersaTune | 15.07 | 4.12 | 10.46 | 3.24 | 17.06 | 8.55 | 10.46 | 22.70 | ||
| 100% Specific Domain | 18.24 | 9.71 | 34.97 | 9.41 | 5.29 | 8.82 | 34.97 | 4.41 | ||
| 2 | Uniform Distribution of Non-Target Domains | 20.59 | 7.94 | 31.08 | 4.71 | 7.69 | 4.98 | 31.08 | 5.23 | |
| VersaTune | 24.70 | 7.35 | 33.92 | 5.02 | 7.08 | 10.58 | 33.92 | 44.69 | ||
| 100% Specific Domain | 23.53 | 9.41 | 55.87 | 17.35 | 14.71 | 7.94 | 55.87 | 25.88 | ||
| 3 | Uniform Distribution of Non-Target Domains | 15.02 | 8.24 | 52.41 | 12.06 | 10.30 | 4.98 | 52.41 | 20.56 | |
| VersaTune | 24.71 | 11.18 | 53.04 | 5.29 | 4.44 | 8.83 | 53.04 | 32.09 | ||
| 100% Specific Domain | 5.00 | 9.12 | 62.89 | 20.29 | 12.94 | 21.76 | 62.89 | 59.11 | ||
| 4 | Uniform Distribution of Non-Target Domains | 5.88 | 5.29 | 56.13 | 14.09 | 13.23 | 20.87 | 56.13 | 47.60 | |
| VersaTune | 14.19 | 8.24 | 58.47 | 13.24 | 8.52 | 9.41 | 58.47 | 25.22 | ||
| Science | 100% Specific Domain | 10.29 | 17.06 | 3.82 | 6.78 | 4.71 | 7.35 | 6.78 | 9.11 | |
| 1 | Uniform Distribution of Non-Target Domains | 6.76 | 17.64 | 9.18 | 5.37 | 7.05 | 4.73 | 5.37 | 13.48 | |
| VersaTune | 3.53 | 17.35 | 7.64 | 8.35 | 16.67 | 6.98 | 8.35 | 29.83 | ||
| 100% Specific Domain | 11.47 | 12.35 | 4.71 | 40.84 | 5.88 | 12.94 | 40.84 | 22.65 | ||
| 2 | Uniform Distribution of Non-Target Domains | 8.24 | 20.59 | 5.68 | 32.78 | 9.12 | 5.85 | 32.78 | 21.30 | |
| VersaTune | 12.36 | 16.17 | 9.43 | 36.97 | 16.89 | 12.65 | 36.97 | 42.78 | ||
| 100% Specific Domain | 21.76 | 7.94 | 8.82 | 63.20 | 4.41 | 10.00 | 63.20 | 37.05 | ||
| 3 | Uniform Distribution of Non-Target Domains | 9.98 | 12.06 | 15.01 | 55.78 | 4.11 | 14.13 | 55.78 | 31.17 | |
| VersaTune | 11.47 | 11.18 | 6.76 | 55.40 | 16.49 | 6.81 | 55.40 | 16.25 | ||
| 100% Specific Domain | 27.94 | 2.35 | 11.47 | 66.15 | 12.65 | 12.59 | 66.15 | 62.30 | ||
| 4 | Uniform Distribution of Non-Target Domains | 13.82 | 13.53 | 11.18 | 58.46 | 6.57 | 6.47 | 58.46 | 24.51 | |
| VersaTune | 10.12 | 10.00 | 4.21 | 61.30 | 5.30 | 6.74 | 61.30 | 16.37 | ||
| Code | 100% Specific Domain | 3.82 | 7.35 | 17.35 | 9.12 | 10.46 | 7.29 | 10.46 | 11.99 | |
| 1 | Uniform Distribution of Non-Target Domains | 3.76 | 5.09 | 7.80 | 7.68 | 5.23 | 5.65 | 5.23 | 6.86 | |
| VersaTune | 5.06 | 11.82 | 8.76 | 8.96 | 5.98 | 8.19 | 5.98 | 15.15 | ||
| 100% Specific Domain | 9.71 | 6.47 | 7.94 | 5.29 | 47.28 | 6.18 | 47.28 | 25.01 | ||
| 2 | Uniform Distribution of Non-Target Domains | 7.90 | 13.49 | 9.05 | 12.03 | 38.31 | 17.76 | 38.31 | 9.19 | |
| VersaTune | 12.22 | 15.04 | 6.78 | 16.28 | 39.77 | 6.14 | 39.77 | 32.02 | ||
| 100% Specific Domain | 22.35 | 8.82 | 14.12 | 7.94 | 61.95 | 10.02 | 61.95 | 47.37 | ||
| 3 | Uniform Distribution of Non-Target Domains | 9.96 | 15.07 | 8.46 | 5.33 | 55.62 | 10.04 | 55.62 | 38.20 | |
| VersaTune | 17.39 | 17.25 | 9.09 | 12.31 | 56.12 | 6.19 | 56.12 | 9.27 | ||
| 100% Specific Domain | 26.18 | 7.06 | 8.82 | 3.24 | 64.76 | 22.65 | 64.76 | 67.95 | ||
| 4 | Uniform Distribution of Non-Target Domains | 14.01 | 13.86 | 6.57 | 6.66 | 58.06 | 13.93 | 58.06 | 41.71 | |
| VersaTune | 5.66 | 4.42 | 10.10 | 9.95 | 59.71 | 13.57 | 59.71 | 14.96 | ||
VersaTune is efficient across different models in both scenarios. Table 4 and Figure 5 show that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 5 and Figure 6, where we averaged the experimental results from LLaMA-3-8B and LLaMA-2-13B. Three interesting phenomena are observed from the outcomes:
Absolute Count vs. Proportion. Notably, by the second epoch of training, the degradation in performance across non-target domains tends to be mitigated to some extent, and in some cases, there is even a positive trend in capability enhancement. We attribute this phenomenon to the fact that the absolute number of samples for each domain, relative to the distribution of domain samples, has a predominant influence on model performance at this stage.
Domain Interactions. Domains are not entirely orthogonal to each other, and there is a degree of mutual reinforcement among them. For instance, enhancing capabilities in the medicine domain can, to a certain degree, boost performance in the science domain, as illustrated in Figure 6.
Domain Mastery Efficiency. From the slope of the target domain performance increase in Figure 6 (b), it is evident that the model’s efficiency in mastering knowledge of a specific domain diminishes over training. In other words, as training progresses, the model’s grasp of the current target domain approaches saturation, while its performance on non-target domains declines sharply. Consequently, greater emphasis should be placed on mitigating losses in non-target domains during this phase, aiming to strike a balance between domain expansion and the salvage of capabilities in non-target domains, which is also demonstrated in Figure 7.
Knowledge consistency training boosts performance. In Table 4 and Figure 5, we also present the experimental results of data composition strategies that allocate domain data in a manner inversely proportional to the pre-existing knowledge distribution detected within each domain. As expected, the inverse distribution strategy yielded even lower performance compared to the simplest approach of uniform distribution, which evenly distributes data across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training, which proves the efficacy of knowledge consistency training stated in Section 3.1.1.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights referring to the knowledge distribution obtained from probing the target model prior to supervised fine-tuning, namely VersaTune Constant, to ablate the component of dynamic adaptation in Algorithm 2. Table 4 and Figure 5 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights throughout the training process by continuously monitoring the learnable potential within each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes during training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial.
[…]